We want to estimate the mean of a population. A random sample of subjects is selected and the sample mean is computed. What is the probability that the sample mean is within 3 units of the true mean if the standard error is 1.8?


1
Expert's answer
2021-10-05T15:51:00-0400

"\mu = ±3 \\
P(\mu-3 < \bar{X} < \mu +3) = P(\frac{(\mu-3) -\mu}{1.8} < \frac{\bar{X} - \mu}{SE} < \frac{(\mu+3)- \mu}{1.8}) \\
= P(-1.67 < Z < 1.67) \\
= P(Z< 1.67) -P(Z < -1.67) \\
= 0.9522 -0.0478 \\
= 0.9044"